OpenStudy (anonymous):

A cylindrical can is to hold 1/2 cup of tomato sauce. What are the dimensions of its diameter and height in centimeters, to minimize the amount of metal to make the can? V=(pi*d^2*h)/4 A=(pi*d^2)+pi*d*h 1qt=946cc

OpenStudy (anonymous):

this is a tricky one....

OpenStudy (amistre64):

are thos answers? or formulas to apply?

OpenStudy (anonymous):

formulas it's a related rates question

OpenStudy (amistre64):

yes, optimizations and such

OpenStudy (anonymous):

yes they want D, H, and Amin

OpenStudy (amistre64):

V = 1/2 cup, soo; .5 = (pi) (r^2) (h) right?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

A = 2 circles and a rectangle :)

OpenStudy (anonymous):

haha :)

OpenStudy (amistre64):

A = 2pi (r^2) + h(2pi r)

OpenStudy (amistre64):

use the volume to calibrate either h or r h = .5/pi r^2 seems easiest right?

OpenStudy (amistre64):

substitute that "value" into the Area formula and then derive

OpenStudy (anonymous):

which value the 946cc?

OpenStudy (amistre64):

the .5 cups...worry about conversions afterwards

OpenStudy (anonymous):

oh okay gotcha

OpenStudy (amistre64):

so...\[h = \frac{1}{2\pi r^2}\] right?

OpenStudy (amistre64):

A = 2pi (r^2) + h(2pi r) sooo \[A = 2\pi r^2 + \frac{2\pi r}{2\pi r^2}\]

OpenStudy (anonymous):

I did get that for A

OpenStudy (amistre64):

\[A = 2\pi r^2 + {1 \over r}\] correct?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

now derive :) you know how to derive?

OpenStudy (anonymous):

yes i do

OpenStudy (amistre64):

A = 4pi r - 1/r^2 then correct?

OpenStudy (amistre64):

A' that is lol

OpenStudy (anonymous):

Yes that is what I just got

OpenStudy (amistre64):

if we combine these with like denoms we get: \[A'=\frac{4\pi r^3 - 1}{r^2}\] correct?

OpenStudy (anonymous):

Yes that looks good

OpenStudy (amistre64):

\[A'=0=4pir^3 -1\] so what are the zeros?

OpenStudy (amistre64):

r = ? the easiest way is to plug this into wolframs site and itll tell us right away ;)

OpenStudy (amistre64):

http://www.wolframalpha.com/

OpenStudy (anonymous):

okay and the equation that I would type in would be \[\sqrt[3]{1/4\pi}\] right

OpenStudy (amistre64):

4pi r^3 - 1 :) and i got abt .4 as the only root

OpenStudy (anonymous):

oh okay you don't have to get r by itself

OpenStudy (amistre64):

not with wolframs site; itll do it for you

OpenStudy (amistre64):

r = aprrox. .430127

OpenStudy (anonymous):

okay and r^2 is the diameter and I can use that to solve for Amin

OpenStudy (amistre64):

Yes; or simply use it to solve 'h'

OpenStudy (anonymous):

Yeah I guess that would be easier lol

OpenStudy (amistre64):

how many cups to a quart? 8 or 4?

OpenStudy (anonymous):

4 cups

OpenStudy (amistre64):

4 cups to 1 quart and we need half a cup so 1/8th of a quart right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

1/2 cup = 946/8 = 118.25 centimeters squared (cc) then

OpenStudy (amistre64):

2 = abt 2.1 centimeters then

OpenStudy (amistre64):

r = abt. 2.1 centimeters

OpenStudy (amistre64):

A = 2pi(2.1)^2 + 1/(2.1)

OpenStudy (amistre64):

Total min area = 28.18 cc

OpenStudy (anonymous):

wouldn't it be 2.1^4 since it was d^2

OpenStudy (anonymous):

oh never mind I see it now

OpenStudy (amistre64):

diameter = 2r = 2(2.1) = 4.2 V = 128.25 = 2pi(2.1)^2 * h 128.25 ---------- = h = abt. 4.63 2pi (2.1)^2

OpenStudy (anonymous):

Thank you very much I've been working on this problem for about 3 hours and you figured it out in 10 minutes lol! You're a lifesaver!

OpenStudy (amistre64):

is it right? cuase i might see an error in it; i allowed the area of the material to be 2 circles, top and bottom of a can; and a height...

OpenStudy (anonymous):

Yes it is it matches up to the parts that I had figured out

OpenStudy (amistre64):

cool :)

OpenStudy (anonymous):

Thanks again!